OpenStudy (nuccioreggie):

mc:radio:(01.02 MC)Tom had $56. He bought a soda for $2 and 2 cookies for $3. Which expression correctly shows the total money Tom has left? A $56 + (−$2) + (−$3) B $56 + (−$2) − (−$3) C $56 − (−$2) − (−$3) D $56 − (−$2) + (−$3) A B C D

OpenStudy (nuccioreggie):

@Littlebird

OpenStudy (nuccioreggie):

i said this one was c

OpenStudy (nuccioreggie):

@Littlebird

OpenStudy (danjs):

Starting money - 1 Soda Price - 2 cookies price = Money Left

OpenStudy (nuccioreggie):

was c correct

OpenStudy (danjs):

no, C is adding the money to his initial 56 dollars 56 - (-2) = 58 Minus a negative number means add

OpenStudy (danjs):

56 - (-2) = 56 + 2 = 58

OpenStudy (danjs):

$56 + (−$2) + (−$3)

OpenStudy (danjs):

Starts with 56, take away 2 dollars for soda, take away 3 dollars for cookies 56 - 2 - 3

OpenStudy (danjs):

plus a negative is the same as subtraction 56 + (-2) = 56 - 2

OpenStudy (danjs):

56 + (-2) + (-3) is what he has left